# SPGCL: Effective Graph Contrastive Learning via SVD-Guided Structural Perturbation

### v1

@@ Abstract
Graph Neural Networks (GNNs) can be highly sensitive to structural noise, including spurious or missing edges caused by adversarial attacks or non-adversarial imperfections. Existing graph contrastive learning methods rely on either random perturbation (e.g., edge dropping) to generate diverse views or purely spectral augmentation (e.g., SVD) to preserve global structural priors. However, random perturbations are structure-agnostic and risk removing critical edges, while purely SVD-based views tend to be dense and lack sufficient diversity. To bridge this gap, we propose SPGCL - a robust graph contrastive learning framework via SVD-guided structural perturbation. Our method couples a lightweight stochastic edge removal with an SVD-guided refinement. This design not only retains the diversity of random augmentations and overcomes the densification issue of spectral methods but also delivers three key advantages: (1) Robustness to accidental deletion: important edges mistakenly removed in the first stage are likely to be recovered by the subsequent SVD-guided step, mitigating harmful effects while preserving diversity. (2) Enrichment with missing links: the SVD-based scoring can introduce new, semantically meaningful edges absent in the original graph, making the contrastive view more informative. (3) Controllable structural discrepancy: by balancing the rates of edge removal and recovery, we explicitly control the edge count between views, ensuring that the contrastive signal stems from semantic structural differences rather than mere edge-number gaps. To further align the two views, we incorporate a contrastive fusion module regularized by a global similarity constraint. Extensive experiments on ten benchmark datasets demonstrate that SPGCL consistently improves the robustness and accuracy of base GNNs, outperforming state-of-the-art graph contrastive learning and structure learning methods.
@@ I Introduction
This dichotomy presents a clear gap: we need a perturbation strategy that is both structure-aware and diversity-preserving. To bridge this gap, we propose SPGCL, a novel framework for robust graph contrastive learning via SVD-guided structural perturbation. Our core innovation is a coupled two-stage augmentation strategy: a lightweight stochastic edge removal followed by an SVD-guided scoring and sparse edge recovery. This design is not merely a combination but a synergistic integration that delivers distinct advantages:
Risk Mitigation for Random Deletion: If the initial random drop accidentally removes an important edge, our SVD-guided recovery step can, with high probability, restore it. This preserves the beneficial diversity introduced by randomness while dramatically reducing its potential harm [12].
@@
We propose a novel SVD-guided structural perturbation strategy for GCL that synergistically couples stochastic edge removal with spectral refinement, addressing the limitations of both random and purely spectral augmentations.
@@ II-B Graph Contrastive Learning and Augmentation Strategies
II-C Beyond Diversity: Towards Rational and Controlled Perturbation
Recent studies have begun to scrutinize the quality of augmentations. Some works point out the instability of random augmentations [31], while others attempt to guide perturbations with node features or gradients [34]. However, these approaches often lack a global structural prior or do not address the post-perturbation recovery of critical edges. Crucially, few methods consider explicitly controlling the structural discrepancy between views (e.g., edge count), leaving open the risk that contrastive learning might exploit trivial, task-irrelevant differences. Our work directly addresses these shortcomings. SPGCL introduces a novel SVD-guided structural perturbation that not only couples random diversity with spectral rationality but also provides a recovery mechanism for accidentally deleted edges, enriches the view with missing high-order links, and enables explicit control over the perturbation scale. This positions SPGCL as a step towards more principled, interpretable, and effective graph augmentation for contrastive learning.
@@ III-B Motivation and Design Intuition
Existing augmentation strategies for graph contrastive learning suffer from a fundamental trade-off: random perturbations (e.g., edge dropping) are diverse but structure-agnostic, while purely spectral augmentations (e.g., SVD) are structure-aware but tend to be dense and deterministic. Our goal is to generate a perturbed view that simultaneously achieves three objectives:
@@
To this end, we propose a two-stage structured perturbation that couples stochastic edge removal with SVD-guided refinement. This design not only balances diversity and structural awareness but also introduces three key advantages:
@@ III-C Two-Stage Structured Perturbation
Stage 1: Lightweight Stochastic Edge Removal.
@@
Stage 2: SVD-Guided Scoring and Sparse Recovery.
@@
To account for the structural impact of the removed edges, we compute a first-order approximation of the singular-value shift ([11, 4]):
A refined score matrix is constructed by adjusting the singular values:
This matrix encodes global structural information while reflecting the perturbation.
@@
where controls the strength of added edges. This yields a sparse, structure-aware augmented view ready for contrastive learning.
@@
Our two-stage perturbation directly addresses the aforementioned objectives:
The random removal ensures diversity.
The SVD-guided scoring preserves global structural patterns and enables recovery of important edges.
The top- sparsification maintains sparsity and allows explicit control over the edge count.
This principled design forms the foundation of SPGCL, enabling robust and informative contrastive learning even under structural noise.
@@ IV The Proposed Method: SPGCL
In this section, we present the overall architecture of SPGCL, which consists of two key components: (1) SVD-guided structural perturbation for generating a sparse, structure-aware augmented view, and (2) contrastive learning-based fusion for aligning the original and perturbed views while enforcing global structural consistency. The complete framework is illustrated in Fig. 1.
@@ V-A1 Datasets
We evaluate SPGCL and baseline models on ten real-world datasets spanning various domains and structural properties. The statistics are summarized in Table I. These include:
@@ V-B Node Classification Performance
V-B1 Overall Comparison
Table II presents the node classification accuracy across ten datasets. We observe:
Traditional GNNs (GCN, SGC, GraphSAGE, GAT) perform well on homophilous graphs (Cora, Citeseer) but degrade significantly on heterophilous ones (Chameleon, Squirrel, Actor).
Graph structure learning methods show improved robustness on some heterophilous datasets (e.g., Pro-GNN on Chameleon), but their performance is inconsistent across graphs.
DropEdge helps alleviate over-smoothing and achieves competitive results on some social networks (BlogCatalog, Flickr) but struggles with complex heterophily.
SPGCL achieves the best or second-best performance on nearly all datasets. It exhibits particularly large gains on challenging heterophilous graphs (e.g., over Pro-GNN on Chameleon, on Squirrel) while maintaining strong accuracy on homophilous ones.
These results validate that our structured perturbation effectively preserves important structural patterns while introducing beneficial diversity, leading to more robust representations.
@@ V-B1 Overall Comparison
V-B2 Comparison with Different Perturbation Strategies
To isolate the benefit of our structured perturbation, we compare SPGCL against two naive perturbation strategies:
Node perturbation: Randomly perturb node features by adding Gaussian noise.
Edge perturbation: Randomly add or drop edges uniformly.
Results in Table V show that SPGCL significantly outperforms both naive strategies across all datasets, with an average improvement of . Edge perturbation generally performs better than node perturbation, confirming that structural noise is a more critical challenge in graph representation learning. The superior performance of SPGCL highlights the importance of coupling stochastic removal with SVD-guided recovery, which prevents arbitrary corruption and enriches the view with high-order connections.
V-C Ablation Study
Ablation on the perturbation module.
We conduct an ablation study to validate the design of our perturbation module, comparing SPGCL with three variants:
@@ V-C Ablation Study
Results in Table III lead to several insights:
SPGCL outperforms all variants, demonstrating the necessity of coupling stochastic and spectral perturbations.
@@
SVD & Edge performs poorly across all datasets. This is because dropping edges after SVD reconstruction randomly destroys the globally consistent patterns recovered by SVD, resulting in a view that is neither diverse nor structure-aware.
These findings underscore that our two-stage design-where stochastic removal provides diversity and SVD-guided recovery preserves and enhances structural integrity-is crucial for generating effective contrastive views.
Ablation on the global consistency loss .
To evaluate the effect of the proposed global similarity-matrix consistency regularizer, we remove from the objective while keeping all other components unchanged. As shown in Table IV, discarding consistently degrades performance on both homophilous (Cora, Citeseer) and heterophilous datasets (Chameleon, Squirrel). This indicates that provides complementary supervision beyond the node-level InfoNCE alignment by stabilizing the global relational structure across the two views.
@@
| Method | Node | Edge | Ours |
| Cora | 68.10 | 75.40 | 83.10 |
| Citeseer | 64.70 | 60.30 | 74.30 |
| CoraML | 85.00 | 85.81 | 90.32 |
| Chameleon | 32.12 | 44.75 | 61.84 |
| Squirrel | 25.47 | 28.81 | 38.62 |
| Wiki | 64.97 | 57.08 | 78.69 |
| BlogCatalog | 62.25 | 66.80 | 79.62 |
| Flickr | 32.35 | 32.00 | 58.04 |
| Amazon-Photo | 82.21 | 82.29 | 90.84 |
| Actor | 24.32 | 24.75 | 30.92 |
V-D Robustness Analysis
To evaluate robustness under structural noise, we simulate increasingly corrupted graphs by randomly removing and of edges, and compare the performance of SPGCL with baselines. Results in Table VI show:
@@ V-D Robustness Analysis
SPGCL maintains the highest or competitive accuracy under both corruption levels, with a smaller relative decline. For example, on Cora, SPGCL drops by only when of edges are removed, whereas GCN drops by .
This resilience stems from our SVD-guided recovery, which can reconstruct important edges even under significant random deletion, and the contrastive fusion, which aligns representations across noisy views.
@@
V-E Visualization of Learned Embeddings

### v2

# SPGCL: Simple yet Powerful Graph Contrastive Learning via SVD-Guided Structural Perturbation

@@ Abstract
Graph Neural Networks (GNNs) are highly sensitive to structural noise, including spurious or missing edges caused by adversarial attacks or non-adversarial imperfections. Existing graph contrastive learning methods typically rely on either random perturbations (e.g., edge dropping) to ensure view diversity, or purely spectral augmentations (e.g., SVD) to preserve global structural priors. However, random perturbations are structure-agnostic and may inadvertently remove critical edges, while purely SVD-based views tend to be dense and lack sufficient diversity. More fundamentally, despite their complementary strengths, there exists no straightforward way to integrate these two fundamentally different paradigms: one operates on discrete edge removal while the other relies on continuous matrix factorization.
To bridge this non-trivial gap, we propose SPGCL, a simple yet powerful framework for robust graph contrastive learning via SVD-guided structural perturbation. Our key insight lies in leveraging a recently developed SVD-based perturbation method, which generalizes the classical structural perturbation theory from symmetric to arbitrary graphs. We creatively adapt this technique to contrastive learning by designing a concise two-stage perturbation strategy: first, a lightweight stochastic edge removal injects necessary diversity; second, truncated SVD is applied to derive a structure-aware scoring matrix, followed by sparse top- edge recovery. This synergistic integration delivers three key advantages: (1) Robustness to accidental deletion - important edges removed in the first stage can be recovered by SVD-guided scoring, mitigating harmful effects while preserving diversity. (2) Enrichment with missing links - the SVD-based scoring can introduce new, semantically meaningful edges absent in the original graph, creating a more informative contrastive view. (3) Controllable structural discrepancy – by balancing edge removal and recovery rates, we explicitly control the edge count between views, ensuring that the contrastive signal stems from semantic structural differences rather than trivial edge-number gaps.
Furthermore, we incorporate a contrastive fusion module regularized by a global similarity constraint to align node embeddings across views. Extensive experiments on ten benchmark datasets demonstrate that SPGCL consistently improves the robustness and accuracy of base GNNs, outperforming state-of-the-art graph contrastive learning and structure learning methods, thereby validating its simplicity, effectiveness, and innovative integration of previously disparate paradigms.
@@ I Introduction
The Fundamental Challenge of Integration. While the complementary strengths of random and spectral augmentations are intuitively appealing, their integration poses a non-trivial challenge. The two paradigms operate on fundamentally different principles: random edge removal is a discrete, structure-agnostic operation, whereas SVD-based augmentation is a continuous, global matrix factorization process. There exists no natural bridge between these two distinct mechanisms, making their synergistic combination far from straightforward. Simply applying both in sequence (e.g., SVD followed by random dropping) often leads to conflicting effects-the global consistency captured by SVD can be arbitrarily destroyed by subsequent random removal. Therefore, despite the clear potential benefits, how to seamlessly and effectively integrate these disparate approaches remains an open and under-explored problem.
Our Insight: SVD-Guided Perturbation as the Bridge. To overcome this challenge, we draw inspiration from a recently developed SVD-based perturbation method [4], which generalizes the classical structural perturbation theory [12] from symmetric to arbitrary graphs via singular value decomposition. This method provides a principled way to quantify the structural impact of edge modifications (addition/removal) through first-order singular value shifts, effectively establishing a continuous link between discrete edge operations and the spectral domain. We recognize that this technique can serve as the missing link to bridge random and spectral augmentations: the stochastic removal step provides the necessary diversity, while the SVD-based perturbation theory enables a structure-aware recovery and enrichment of edges, ensuring global consistency is preserved.
Proposed Framework: SPGCL. Building on this insight, we propose SPGCL, a simple yet powerful framework for robust graph contrastive learning via SVD-guided structural perturbation. Our core innovation is a coupled two-stage augmentation strategy: a lightweight stochastic edge removal followed by an SVD-guided scoring and sparse edge recovery. This design is not merely a combination but a synergistic integration that delivers distinct advantages:
Risk Mitigation for Random Deletion: If the initial random drop accidentally removes an important edge, our SVD-guided recovery step can, with high probability, restore it. This preserves the beneficial diversity introduced by randomness while dramatically reducing its potential harm.
@@
We identify and articulate the fundamental challenge of integrating random and spectral augmentations in graph contrastive learning, highlighting the disparate nature of the two paradigms.
We introduce a novel SVD-guided structural perturbation strategy that creatively adapts a recent SVD-based perturbation method [4] to bridge random diversity with spectral rationality, enabling their synergistic integration for the first time.
@@ II-B Graph Contrastive Learning and Augmentation Strategies
II-C Structural Perturbation and Its SVD-based Generalization
Structural Perturbation Method (SPM) pioneered by Lü et al. [12] provides a principled framework for link prediction by analyzing the stability of a network’s eigenvectors under small perturbations. Rooted in first-order eigenvalue perturbation theory, SPM offers a global, structure-aware perspective on link predictability, moving beyond local heuristics. However, it is inherently limited to symmetric, undirected graphs due to its reliance on eigendecomposition, restricting its direct application to general graph types.
To overcome this limitation, GADPN [4] generalizes SPM to arbitrary (including directed and attributed) graphs via Singular Value Decomposition (SVD). By deriving a first-order singular value perturbation formula , GADPN extends the original eigenvalue-based perturbation theory to general adjacency matrices. This extension not only enables structure-aware edge recovery and denoising in broader graph learning scenarios but also establishes a continuous link between discrete edge modifications and the spectral domain-a crucial insight that bridges operations on graph structures with their low-rank approximations.
While GADPN focuses on graph structure learning for semi-supervised node classification, its core SVD-based perturbation technique provides a generic and powerful tool to quantify and compensate for structural changes. In this work, we recognize that this very technique can be creatively repurposed to address a fundamentally different challenge: integrating random and spectral augmentations in graph contrastive learning. SPGCL thus stands as an innovative application of the generalized perturbation framework, demonstrating its versatility and potential to connect previously disparate paradigms in graph representation learning.
II-D Beyond Diversity: Towards Rational and Controlled Perturbation
Recent studies have begun to scrutinize the quality of augmentations. Some works point out the instability of random augmentations [31], while others attempt to guide perturbations with node features or gradients [34]. However, these approaches often lack a global structural prior or do not address the post-perturbation recovery of critical edges. Crucially, few methods consider explicitly controlling the structural discrepancy between views (e.g., edge count), leaving open the risk that contrastive learning might exploit trivial, task-irrelevant differences.
Our work directly addresses these shortcomings. SPGCL introduces a novel SVD-guided structural perturbation that not only couples random diversity with spectral rationality but also provides a recovery mechanism for accidentally deleted edges, enriches the view with missing high-order links, and enables explicit control over the perturbation scale. By building upon the generalized perturbation theory reviewed in Section II-C, our method offers a principled and interpretable approach to graph augmentation, positioning SPGCL as a step towards more effective and robust contrastive learning on graphs.
@@ III-B Motivation and Design Intuition
Existing augmentation strategies for graph contrastive learning suffer from a fundamental trade-off: random perturbations (e.g., edge dropping) are diverse but structure-agnostic, while purely spectral augmentations (e.g., SVD) are structure-aware but tend to be dense and deterministic. To overcome this trade-off and achieve robust representations, our goal is to generate a perturbed view that simultaneously fulfills three objectives:
@@
These objectives directly motivate the three key advantages of our method: robustness to accidental deletion, enrichment with missing links, and controllable structural discrepancy.
To this end, we propose a two-stage structured perturbation that couples stochastic edge removal with SVD-guided refinement. This simple yet principled design not only balances diversity and structural awareness but also naturally delivers the aforementioned advantages:
@@ III-C Two-Stage Structured Perturbation
Stage 1: Lightweight Stochastic Edge Removal
@@
Stage 2: SVD-Guided Scoring and Sparse Recovery
@@
In practice, we adopt the randomized SVD algorithm [8] to efficiently compute the rank- approximation.
Key step: estimating structural impact via generalized perturbation. To account for the structural impact of the removed edges in a principled manner, we leverage the generalized singular value perturbation theory recently developed in [4]. This theory extends the classic eigenvalue perturbation method [12] from symmetric to arbitrary graphs via SVD, providing a first-order approximation of how edge modifications affect the singular values:
Eq. (5) is crucial: it establishes a continuous, differentiable relationship between discrete edge removal () and the spectral domain (). This exactly serves as the missing bridge that allows us to connect the stochastic removal in Stage 1 with the spectral recovery in Stage 2.
A refined score matrix is then constructed by adjusting the singular values according to their estimated shifts:
This matrix encodes global structural information while explicitly reflecting the perturbation introduced by the removed edges.
@@
where controls the strength (weight) of the added edges. This yields a sparse, structure-aware augmented view ready for contrastive learning.
@@
Our two-stage perturbation directly addresses the objectives outlined in Sec. III-B and embodies the simple yet powerful design philosophy:
The random removal ensures diversity in a lightweight manner.
The SVD-guided scoring, built upon the generalized perturbation theory [4], preserves global structural patterns and enables recovery of important edges by mathematically linking edge-level changes to spectral shifts.
The top- sparsification maintains sparsity and provides explicit control over the edge count and structural discrepancy.
This principled and modular design, grounded in a rigorous perturbation framework, forms the foundation of SPGCL, enabling robust and informative contrastive learning even under structural noise with minimal complexity overhead.
@@ IV The Proposed Method: SPGCL
In this section, we present the overall architecture of SPGCL, a simple yet powerful framework for robust graph contrastive learning. It consists of two seamlessly integrated components: (1) an SVD-guided structural perturbation module that generates a sparse, structure-aware augmented view, and (2) a contrastive learning-based fusion module that aligns the original and perturbed views while enforcing global structural consistency. The complete pipeline is illustrated in Fig. 1.
@@ IV-B Contrastive Learning-Based Fusion
Summary
The proposed SPGCL framework elegantly combines a principled two-stage perturbation with a multi-level contrastive fusion. The perturbation module ensures that the augmented view is both diverse and structure-aware, while the fusion module aligns the two views at node and global levels to learn robust and consistent representations. This integrated design achieves a favorable balance between simplicity and effectiveness, as evidenced by the extensive experiments in the following section.
@@ V-A1 Datasets
We evaluate SPGCL and baseline models on ten real-world datasets spanning various domains and structural properties, with statistics summarized in Table I. These include:
@@ V-B Node Classification Performance
V-B1 Comparison with Baselines and Perturbation Strategies
The overall node classification accuracy across all datasets is presented in Table II. We also include a comparison against two naive perturbation strategies—Node perturbation (adding Gaussian noise to features) and Edge perturbation (random edge addition/dropping)—in Table III to isolate the benefit of our structured design.
Key observations are as follows:
Traditional GNNs (GCN, SGC, GraphSAGE, GAT) perform well on homophilous graphs (Cora, Citeseer) but degrade significantly on heterophilous ones (Chameleon, Squirrel, Actor), highlighting their sensitivity to structural mismatch.
Graph structure learning methods (LDS, Pro-GNN) show improved robustness on some heterophilous datasets (e.g., Pro-GNN on Chameleon), but their performance is inconsistent across different graph types.
DropEdge helps alleviate over-smoothing on homophilous graphs and achieves competitive results on BlogCatalog and Flickr, but still fails to generalize well under complex structural heterophily.
SPGCL consistently achieves the best or second-best performance on nearly all datasets. It exhibits particularly large gains on challenging heterophilous graphs (e.g., +5.6% over Pro-GNN on Chameleon, +3.6% on Squirrel) while maintaining strong accuracy on homophilous ones. This demonstrates the broad effectiveness of our structured perturbation.
SPGCL significantly outperforms naive perturbation strategies (Table III), with an average improvement of 10-15%. This confirms that the benefits stem from our principled, structure-aware design rather than from random noise injection.
These results validate that SPGCL’s coupled stochastic-SVD perturbation, built upon the generalized perturbation framework [4], effectively preserves important global structural patterns while introducing beneficial local diversity. The consistent gains across both homophilous and heterophilous graphs demonstrate that our method successfully bridges the long-standing gap between random and spectral augmentations, leading to more robust and discriminative node representations.
@@ V-B1 Comparison with Baselines and Perturbation Strategies
| Method | Node | Edge | Ours |
| Cora | 68.10 | 75.40 | 83.10 |
| Citeseer | 64.70 | 60.30 | 74.30 |
| CoraML | 85.00 | 85.81 | 90.32 |
| Chameleon | 32.12 | 44.75 | 61.84 |
| Squirrel | 25.47 | 28.81 | 38.62 |
| Wiki | 64.97 | 57.08 | 78.69 |
| BlogCatalog | 62.25 | 66.80 | 79.62 |
| Flickr | 32.35 | 32.00 | 58.04 |
| Amazon-Photo | 82.21 | 82.29 | 90.84 |
| Actor | 24.32 | 24.75 | 30.92 |
V-C Ablation Studies
V-C1 Ablation on the Perturbation Module
To validate the design of our perturbation module, we compare SPGCL with three variants:
@@ V-C1 Ablation on the Perturbation Module
Results in Table IV lead to several insights:
SPGCL outperforms all variants, demonstrating the necessity of coupling stochastic and spectral perturbations. Neither component alone is sufficient.
@@
SVD & Edge performs poorly across all datasets. This is because dropping edges after SVD reconstruction randomly destroys the globally consistent patterns recovered by SVD, resulting in a view that is neither diverse nor structure-aware, highlighting the importance of our specific integration order.
These findings underscore that our two-stage design-where stochastic removal provides diversity and SVD-guided recovery, grounded in the perturbation theory of [4], preserves and enhances structural integrity-is crucial for generating effective contrastive views. The poor performance of the “SVD & Edge” variant further confirms that simply cascading the two operations fails; the principled integration enabled by the perturbation framework is key.
@@
V-C2 Ablation on the Global Consistency Loss
To evaluate the contribution of the proposed global similarity-matrix consistency regularizer, we remove from the objective while keeping all other components unchanged. As shown in Table V, discarding consistently leads to a performance drop on both homophilous (Cora, Citeseer) and heterophilous (Chameleon, Squirrel) datasets. This indicates that provides complementary supervision beyond the node-level InfoNCE alignment by explicitly stabilizing the global relational structure across the two views, which is particularly beneficial for maintaining structural consistency under perturbation.
@@ V-C2 Ablation on the Global Consistency Loss
V-D Robustness Analysis under Structural Noise
To evaluate robustness under increasing structural noise, we simulate corrupted graphs by randomly removing 20% and 30% of edges, and compare SPGCL with baselines. Results in Table VI show:
@@ V-D Robustness Analysis under Structural Noise
SPGCL maintains the highest or competitive accuracy under both corruption levels, with a smaller relative decline. For example, on Cora, SPGCL drops by only 6.4% when 30% of edges are removed, whereas GCN drops by 9.7%.
This resilience stems from the theoretically grounded recovery mechanism provided by the generalized perturbation framework [4], which can reconstruct important edges even under significant random deletion, coupled with the contrastive fusion that aligns representations across noisy views. This demonstrates SPGCL’s practical value in noisy real-world scenarios and further validates the utility of the underlying perturbation theory beyond its original domain.
@@
V-E Parameter Sensitivity Analysis
We analyze the sensitivity of SPGCL to two key hyperparameters: the base edge drop ratio and the offset in Eq. (7), which jointly control the scale of perturbation (). We perform a 2D grid search over and and report the resulting accuracy on Cora and Squirrel in Fig. 2.
Observations: The performance exhibits a clear unimodal trend. When is too small (under-perturbation), the contrastive view is overly similar to the original, limiting diversity. As increases to a moderate range, performance improves and reaches a peak. However, overly large (over-perturbation) introduces many low-confidence edges, leading to graph densification and performance degradation. This analysis confirms that our method is not overly sensitive to hyperparameters within a reasonable range and provides practical guidance for tuning.
V-F Visualization of Learned Embeddings